2. For Students. An electron is released in a uniform electric field and experiences an electric force of 1.6 x 10 N toward the south. Find the magnitude and direction of the electric field.
Answer #1

Charge on electron e = - 1.6*10^(-19) C

Electric Force = q*E

Magnitude of E = F/q = 1.6*10^(-14)/(1.6*10^(-19)) = 10^5 N/C

Direction should be opposite to electric force so

Direction is toward the North.